# October   ||   2024 AI-Driven Control of Bioelectric signalling

### v1

@@ 1.1 Bioelectric Signaling in Morphogenesis and Regeneration
McLaughlin and Levin [McLaughlin2018] explored bioelectric signaling in regeneration, demonstrating that specific voltage modulations can direct tissue structure outcomes. Their experiments with Xenopus tadpoles involved manipulating transmembrane potentials (Vmem) using ion channel overexpression and optogenetic tools to induce or inhibit limb regeneration. Pharmacological agents such as ouabain were used to influence ion flux, while optogenetic tools allowed precise, spatially resolved control over bioelectric states through light-induced activation of specific ion channels. Live imaging tracked morphological changes, and genetic markers identified differentiation in response to bioelectric alterations. Results showed that targeted Vmem modulations led to significant changes in regeneration patterns, promoting or inhibiting cell proliferation and differentiation as needed. For example, hyperpolarising cells promoted directed differentiation, while depolarizing states increased proliferation. These findings support the concept of a bioelectric code as an instructive signal for tissue organization and regenerative processes, further validating the therapeutic potential of bioelectric modulation in regenerative medicine by providing positional information crucial for proper morphogenesis.
@@
Practical applications of bioelectric principles are further exemplified by Houpu Li et al. [Li2023], who design, fabricate, and test a wearable bioelectronic bandage capable of delivering fluoxetine to wounds to enhance healing. The methodology involved designing a bioelectronic bandage comprising an ion pump delivery module and a controller module. The ion pump, fabricated using polydimethylsiloxane (PDMS), utilized hydrogel-filled capillaries as ion-selective membranes to allow positively charged fluoxetine ions to pass while blocking other ions. The controller module included a programmable printed circuit board (PCB) managing precise fluoxetine dosage through electrophoretic force. In vivo testing on C57B6 male mice demonstrated that fluoxetine delivery from the bioelectronic bandage significantly increased re-epithelialization by 39.9% and reduced the M1/M2 macrophage ratio by 27.2%, indicating enhanced wound progression and reduced inflammation. These results illustrate the therapeutic potential of programmable bioelectric devices in medical applications, showcasing the integration of bioelectric principles with wearable technology for targeted drug delivery and improved healing outcomes.
@@
Further supporting the role of bioelectric signaling in regeneration, Beane et al. [Beane2013] investigate how H,K-ATPase ion transporters influence tissue remodeling and organ scaling in planarian regeneration. Their experiments with Schmidtea mediterranea planarians involved RNA interference (RNAi) to inhibit H,K-ATPase function, resulting in regenerates with disproportionately large pharynges and shrunken heads. Apoptosis assays revealed that H,K-ATPase inhibition led to reduced apoptosis during the later stages of regeneration, disrupting normal tissue remodeling and causing pre-existing tissues to retain their original size and shape instead of adjusting proportionally to the regenerating body. Membrane potential measurements and intracellular pH recordings confirmed changes in cellular ion states due to H,K-ATPase RNAi. Immunofluorescence and in situ hybridization techniques further demonstrated disruptions in tissue remodeling markers, underscoring the essential role of bioelectric regulation in maintaining tissue homeostasis and proper organ scaling during regenerative processes.
@@
In the realm of collective behavior and stress management, Shreesha and Levin [SHREESHA2024150396] present a computational model where stress sharing among cells facilitates coordinated morphogenesis. The model simulates a virtual embryo on a two-dimensional grid, where each cell can exhibit one of two states representing different cell types. Cells are randomly arranged initially, forming a scrambled configuration that deviates from the target pattern. Each cell experiences stress quantified as a binary signal indicating whether it is in its correct target position. In stress-sharing configurations, stressed cells can diffuse their stress to neighbors, creating a shared stress environment that encourages collective movement toward target positions. Genetic algorithms evolved these virtual embryos over generations, assessing fitness based on the L2 distance from the target pattern. Results showed that stress-sharing embryos achieved higher fitness faster than those without stress-sharing or with hardwired configurations. Stress-sharing increased the movement range for individual cells, allowing them to resolve their positions more effectively and achieve coherent tissue organization.
@@
Hansali et al. [Hansali2024] examine the influence of bioelectrical patterns on morphogenesis through evolutionary simulations and empirical validation in planarian regeneration. Utilizing a neural cellular automata (NCA) model, the study investigates how different bioelectric encodings - direct, indirect, and binary triggers - contribute to developmental outcomes. The methodology involved simulating 9x9 cell grids representing artificial organisms, each controlled by a neural network optimized via a genetic algorithm over 250 generations. Target morphologies, such as the "Tadpole" and the "French Flag," inspired by actual biological morphogenetic processes, were used to assess the effectiveness of different encoding strategies. Results demonstrated that direct pattern organisms reliably achieved target morphologies even after bioelectric resetting due to their linear encoding. Indirect pattern organisms exhibited emergent robustness against partial bioelectric disruptions but struggled with complete resets, while binary trigger organisms were highly sensitive to precise timing of bioelectric cues, failing morphogenesis under altered conditions. Empirical validation through SSRI experiments in planarian regeneration confirmed that SSRIs disrupt bioelectric interpretation, resulting in variable morphologies and bistable anatomical states. This underscores the critical role of bioelectric signaling in developmental regulation and the robustness of indirect encoding strategies against environmental perturbations.
@@ 1.2 Synthetic Biology and Multicellularity
The challenges in synthetic multicellularity are comprehensively addressed by [Sole2024], who categorize synthetic multicellular systems into circuits, programmable assemblies, and synthetic morphologies. The methodology spans genetic modification to create complex cellular responses, such as Boolean logic gates and pattern formation circuits, as well as manipulating cell adhesion properties to influence spatial organization within cell collectives. Experiments in synthetic multicellular circuits involved designing multicellular consortia where cellular strains communicate through chemical signals in liquid media or agar plates, resulting in stable configurations or patterns under controlled conditions. Programmable synthetic assemblies utilized differential adhesion models to drive self-organized cellular aggregates, achieving predictable spatial arrangements based on specific initial conditions. Synthetic morphology experiments, including the development of Xenobots and Anthrobots, emphasized the emergence of functionalities like memory and homeostasis without genetic editing, showcasing the balance between engineered predictability and emergent biological behaviors. The results highlight both the successes and inherent challenges in scaling synthetic multicellular systems, pointing to the necessity for novel methodologies to manage complexity and unpredictability in larger-scale constructs.
@@
Advancements in bioinformatics and network analysis have further elucidated the connections between bioelectric signaling and disease. Pio-Lopez and Levin [PioLopez2024] introduce MultiXVERSE, a universal multilayer network embedding method applied to investigate the causal link between GABA neurotransmitters and cancer. By constructing a multiplex-heterogeneous network that combines drug, disease, and gene interactions, MultiXVERSE employs Random Walks with Restart (RWR) for node similarity calculations and link prediction via a Random Forest classifier. The study involved creating a multilayer network incorporating protein-protein interactions, drug interactions from sources like DrugBank, and disease networks representing symptom similarity and comorbidity relationships. Experimental validation using Xenopus laevis embryos exposed to the GABA agonist muscimol demonstrated that fluoxetine delivery via a wearable bioelectronic bandage significantly enhanced wound healing, evidenced by a 39.9% increase in re-epithelialization and a 27.2% reduction in the M1/M2 macrophage ratio. These findings suggest that bioelectric states, modulated by neurotransmitters like GABA, play a significant role in tumorigenesis and cancer progression, offering potential avenues for therapeutic interventions through bioelectric modulation.
@@
Exploring the evolutionary aspects of multicellular intelligence, Hartl, Risi, and Levin [Hartl2024] investigate how different levels of cellular competency within a multi-scale competency architecture (MCA) affect the evolutionary process. Their in-silico model simulates morphogenesis using neural cellular automata (NCAs), where cellular agents evolve to form target patterns. The methodology involves setting up NCAs on an 8x8 grid to form predefined patterns, such as a Czech flag, through coordinated cell state updates. Each cell contains an artificial neural network (ANN) that processes information from neighboring cells and proposes state changes. Key competency parameters, including decision-making probability (PD) and redundancy in decision pathways (R), are adjusted to simulate various levels of cellular reliability and computational capacity. The evolutionary fitness of each NCA is measured based on the accuracy and stability of forming the target pattern, with cells experiencing noise to simulate environmental challenges. Results indicate that higher competency levels accelerate evolutionary success and enhance the robustness and adaptability of morphogenetic processes. Additionally, NCAs with multi-scale competency encoding demonstrated better generalization to new environmental conditions compared to direct encoding models, suggesting that cellular-level intelligence significantly contributes to the adaptability and resilience of developing structures.
@@
The concept of memory as a dynamic and adaptive process is further explored by Levin [Levin2024], who posits that biological memories are not static repositories but are continuously reinterpreted to maintain functional relevance during significant physiological changes. Through conceptual analysis and comparative studies, Levin illustrates how memory engrams adapt to new biological contexts, such as metamorphosis in caterpillars and planarian regeneration. For instance, during metamorphosis, memories associated with a caterpillar’s sensory-motor systems are reformatted to remain functionally relevant for the butterfly’s needs, despite drastic physical and behavioral transformations. Similarly, planaria can transfer memories through body regeneration, revealing that specific memory cues adapt across different biological states. This phenomenon, termed "mnemonic improvisation," is central to understanding intelligence, as memory supports flexible, real-time problem-solving and continuity of self. Levin’s perspective offers a framework for understanding intelligence as an emergent property of dynamic memory processes that evolve across different biological scales.
@@
Investigating the emergence of higher-level agency, Tissot et al. [Tissot2024] examine how phase synchronization among individual decision cycles in a decentralized system can lead to the emergence of higher-level agency. Their computational model simulates collective behaviors of individual units that engage in synchronized decision-making, achieving forms of collective action that surpass the capabilities of individual units acting alone. Each unit operates within a modular system, interacting more frequently with intra-module units than with inter-module units. The decision cycles are governed by oscillatory functions where phase adjustments occur based on the alignment of decisions within each module. Results demonstrate that synchronization leads to spontaneous collective shifts from local to global optima, overcoming individual-level stress barriers. After synchronization, modules coordinate responses across the entire system, achieving states with minimized collective stress. These findings suggest that synchronization mechanisms are fundamental to the emergence of collective intelligence in decentralized biological systems, enabling higher-level problem-solving competencies.
@@
In cancer research, Kofman and Levin [Kofman2024] perform a meta-analysis evaluating pharmacological agents targeting ion channels as potential cancer therapies. Their comprehensive literature search across databases like IUPHAR, DrugBank, and PubMed identified various ion channel-targeting drugs, categorizing them based on their cancer-promoting or inhibiting effects. The analysis revealed that numerous ion channel drugs exhibit significant cancer-modifying properties across multiple cancer types, including breast, liver, brain, and ovarian cancers. Voltage-gated sodium and potassium channels frequently appeared in both pro- and anti-cancer categories, suggesting that the bioelectric state of cells can significantly influence tumor progression and metastasis. Some drugs demonstrated dual effects, indicating that factors such as timing, dosage, or biological context might influence their impact on cancer phenotypes. These findings highlight the therapeutic potential of bioelectric modulation in cancer treatment strategies, offering new targets for intervention beyond traditional genetic approaches.
@@
Zhang, Goldstein, and Levin [zhang2023classicalsortingalgorithmsmodel] explore morphogenetic processes through the application of classical sorting algorithms. By implementing sorting algorithms such as Bubble Sort, Insertion Sort, and Selection Sort in a decentralized, cell-view model, they simulate self-sorting cellular arrays capable of handling error-prone conditions through emergent behaviors like clustering and delayed gratification. The methodology involved treating each element in an array as an autonomous agent capable of viewing its neighboring elements and deciding whether to swap based on local conditions. In addition to normal, fully-functional cells, the study introduced "frozen" cells that could either initiate no swaps or be immovable by others, simulating unreliability as seen in biological media. Results demonstrated that cell-view algorithms maintained sorting efficiency and robustness even in the presence of unreliable cells, showcasing emergent problem-solving behaviors in minimalistic systems. These findings indicate that algorithmic models can effectively mimic biological self-organization and problem-solving capabilities, offering insights into the decentralized control mechanisms inherent in biological morphogenesis.
@@
Theoretical frameworks for self-organization during prenatal development are advanced by Ciaunica et al. [Ciaunica2023], who apply the concept of Markov blankets to model the maternal-fetal relationship during human prenatal development. Their study emphasizes the interplay between maternal and fetal immune systems, conceptualizing pregnancy as a dynamic process where two immune systems operate as self-organizing entities with distinct yet interconnected Markov blankets. The placenta serves as a boundary regulating nutrient and signal exchange, maintaining separate yet coordinated environments for both the mother and fetus. The methodology involved a detailed analysis of immune interactions, including the migration of immune cells like macrophages and natural killer cells through the placenta, contributing to the unique immune environment of pregnancy. The results highlight a theoretical framework where the Markov blanket formalism provides insights into the maternal-fetal relationship as an instance of nested selves, facilitating co-homeostasis and coordinated developmental processes. This perspective offers a novel understanding of biological self-organization, emphasizing the role of shared Markov blankets in maintaining the balance and stability of interconnected yet distinct biological systems during prenatal development.
@@
Decentralized control mechanisms in microswimmers are investigated by Hartl, Levin, and Zöttl [Hartl], who demonstrate that artificial neural networks (ANNs) and genetic algorithms can optimize collective locomotion strategies in N-bead swimmers modeled after the Najafi-Golestanian (NG) microswimmer. The methodology involved constructing a virtual N-bead swimmer model where each bead acts as an independent agent capable of perceiving the states of its neighboring beads and applying forces accordingly. Using reinforcement learning techniques, the ANN parameters were evolved to maximize the swimmer’s speed, with the study testing swimmer sizes ranging from three to 100 beads. Results showed that the decentralized approach enabled robust and efficient swimming gaits, with locomotion strategies scaling effectively as the number of beads increased. Type B swimmers, in particular, displayed collective contractions reminiscent of caterpillar-like movement. Additionally, the evolved locomotion strategies demonstrated resilience to morphological changes and adaptability to different swimmer configurations without retraining, highlighting the potential of decentralized systems in achieving coordinated movement in autonomous microswimmers for biomedical applications.
@@
Levin [Michael2019] theorizes that bioelectric networks in multicellular systems establish cognitive boundaries that enable cells to function as unified cognitive entities. By integrating developmental biology, bioelectric signaling, and cognitive science, Levin posits that bioelectric gradients serve as computational mediums guiding multicellular coordination and pattern formation. The methodology involves a theoretical framework combining bioelectric signaling mechanisms, such as ion channels and gap junctions, with cognitive boundary concepts to model how cells collectively process information and regulate behavior. Examples from studies on Xenopus embryos, planarians, and tadpole regeneration support the hypothesis that bioelectric states function as pre-patterns, directing cellular organization and overriding genetic instructions to achieve specific anatomical outcomes. The results indicate that bioelectric signaling underpins cognitive-like processes across biological scales, suggesting that the "self" emerges from bioelectric boundaries that define the spatio-temporal limits within which cells interact and maintain coherence as a unified system.
@@
Lyon et al. [Lyon2021] expand the understanding of cognitive processes in biology by redefining basal cognition to include sensory and information-processing mechanisms observed across diverse life forms. Their study examines cognitive behaviors in single-celled organisms, prokaryotes, and plants, focusing on capacities such as sensing, perception, memory, decision-making, and communication. By investigating mechanisms like quorum sensing in bacteria, cellular adhesion in amoebas, and bioelectrical signaling in plants, the authors argue that foundational cognitive capacities are evolutionarily conserved and precede the development of nervous systems. Experiments involved monitoring cellular responses to nutrient gradients and light stimuli, eliciting directional movement and adaptive behaviors without the presence of a nervous system. The results reveal that even simple organisms exhibit basic cognitive functions, suggesting a continuity of cognitive processes across biological complexity. This foundational understanding informs the study of more advanced cognitive systems and has potential applications in regenerative medicine and artificial intelligence, where principles of basal cognition can inspire novel approaches to problem-solving and adaptive behaviors.
@@
In "Resting Potential, Oncogene-induced Tumorigenesis, and Metastasis: The Bioelectric Basis of Cancer in vivo," Lobikin et al. [Lobikin2012] utilize Xenopus laevis tadpoles to investigate how bioelectric signaling influences cancerous transformations. The methodology involved modulating transmembrane voltage potentials (Vmem) by depolarizing specific GlyCl-expressing instructor cells using pharmacological agents like ivermectin to open glycine-gated chloride channels. By manipulating extracellular chloride concentrations, controlled depolarization was induced, leading to melanocytes exhibiting metastatic-like transformations, including hyperproliferation, morphological changes, and invasive behavior. Additionally, exposure to the carcinogen 4-Nitroquinoline 1-oxide (4NQO) generated localized tumors, with bioelectric changes visualized using fluorescent voltage- and ion-reporter dyes. The study found that oncogene-induced tumors displayed elevated sodium levels, potentially useful for non-invasive diagnostics. Conversely, forced hyperpolarization of cells via expression of hyperpolarizing ion channels mitigated oncogene-induced tumorigenesis, demonstrating that bioelectric state modulation can suppress tumor formation. These findings highlight the active role of bioelectric signaling in cancer progression and metastasis, suggesting bioelectric modulation as a promising strategy for cancer therapeutics.

### v2

@@ 1.1 Bioelectric Signaling in Morphogenesis and Regeneration
McLaughlin and Levin [28] explored bioelectric signaling in regeneration, demonstrating that specific voltage modulations can direct tissue structure outcomes. Their experiments with Xenopus tadpoles involved manipulating transmembrane potentials (Vmem) using ion channel overexpression and optogenetic tools to induce or inhibit limb regeneration. Pharmacological agents such as ouabain were used to influence ion flux, while optogenetic tools allowed precise, spatially resolved control over bioelectric states through light-induced activation of specific ion channels. Live imaging tracked morphological changes, and genetic markers identified differentiation in response to bioelectric alterations. Results showed that targeted Vmem modulations led to significant changes in regeneration patterns, promoting or inhibiting cell proliferation and differentiation as needed. For example, hyperpolarising cells promoted directed differentiation, while depolarizing states increased proliferation. These findings support the concept of a bioelectric code as an instructive signal for tissue organization and regenerative processes, further validating the therapeutic potential of bioelectric modulation in regenerative medicine by providing positional information crucial for proper morphogenesis.
@@
Practical applications of bioelectric principles are further exemplified by Houpu Li et al. [23], who design, fabricate, and test a wearable bioelectronic bandage capable of delivering fluoxetine to wounds to enhance healing. The methodology involved designing a bioelectronic bandage comprising an ion pump delivery module and a controller module. The ion pump, fabricated using polydimethylsiloxane (PDMS), utilized hydrogel-filled capillaries as ion-selective membranes to allow positively charged fluoxetine ions to pass while blocking other ions. The controller module included a programmable printed circuit board (PCB) managing precise fluoxetine dosage through electrophoretic force. In vivo testing on C57B6 male mice demonstrated that fluoxetine delivery from the bioelectronic bandage significantly increased re-epithelialization by 39.9% and reduced the M1/M2 macrophage ratio by 27.2%, indicating enhanced wound progression and reduced inflammation. These results illustrate the therapeutic potential of programmable bioelectric devices in medical applications, showcasing the integration of bioelectric principles with wearable technology for targeted drug delivery and improved healing outcomes.
@@
Further supporting the role of bioelectric signaling in regeneration, Beane et al. [2] investigate how H,K-ATPase ion transporters influence tissue remodeling and organ scaling in planarian regeneration. Their experiments with Schmidtea mediterranea planarians involved RNA interference (RNAi) to inhibit H,K-ATPase function, resulting in regenerates with disproportionately large pharynges and shrunken heads. Apoptosis assays revealed that H,K-ATPase inhibition led to reduced apoptosis during the later stages of regeneration, disrupting normal tissue remodeling and causing pre-existing tissues to retain their original size and shape instead of adjusting proportionally to the regenerating body. Membrane potential measurements and intracellular pH recordings confirmed changes in cellular ion states due to H,K-ATPase RNAi. Immunofluorescence and in situ hybridization techniques further demonstrated disruptions in tissue remodeling markers, underscoring the essential role of bioelectric regulation in maintaining tissue homeostasis and proper organ scaling during regenerative processes.
@@
In the realm of collective behavior and stress management, Shreesha and Levin [39] present a computational model where stress sharing among cells facilitates coordinated morphogenesis. The model simulates a virtual embryo on a two-dimensional grid, where each cell can exhibit one of two states representing different cell types. Cells are randomly arranged initially, forming a scrambled configuration that deviates from the target pattern. Each cell experiences stress quantified as a binary signal indicating whether it is in its correct target position. In stress-sharing configurations, stressed cells can diffuse their stress to neighbors, creating a shared stress environment that encourages collective movement toward target positions. Genetic algorithms evolved these virtual embryos over generations, assessing fitness based on the L2 distance from the target pattern. Results showed that stress-sharing embryos achieved higher fitness faster than those without stress-sharing or with hardwired configurations. Stress-sharing increased the movement range for individual cells, allowing them to resolve their positions more effectively and achieve coherent tissue organization.
@@
Hansali et al. [12] examine the influence of bioelectrical patterns on morphogenesis through evolutionary simulations and empirical validation in planarian regeneration. Utilizing a neural cellular automata (NCA) model, the study investigates how different bioelectric encodings - direct, indirect, and binary triggers - contribute to developmental outcomes. The methodology involved simulating 9x9 cell grids representing artificial organisms, each controlled by a neural network optimized via a genetic algorithm over 250 generations. Target morphologies, such as the "Tadpole" and the "French Flag," inspired by actual biological morphogenetic processes, were used to assess the effectiveness of different encoding strategies. Results demonstrated that direct pattern organisms reliably achieved target morphologies even after bioelectric resetting due to their linear encoding. Indirect pattern organisms exhibited emergent robustness against partial bioelectric disruptions but struggled with complete resets, while binary trigger organisms were highly sensitive to precise timing of bioelectric cues, failing morphogenesis under altered conditions. Empirical validation through SSRI experiments in planarian regeneration confirmed that SSRIs disrupt bioelectric interpretation, resulting in variable morphologies and bistable anatomical states. This underscores the critical role of bioelectric signaling in developmental regulation and the robustness of indirect encoding strategies against environmental perturbations.
@@ 1.2 Synthetic Biology and Multicellularity
The challenges in synthetic multicellularity are comprehensively addressed by [42], who categorize synthetic multicellular systems into circuits, programmable assemblies, and synthetic morphologies. The methodology spans genetic modification to create complex cellular responses, such as Boolean logic gates and pattern formation circuits, as well as manipulating cell adhesion properties to influence spatial organization within cell collectives. Experiments in synthetic multicellular circuits involved designing multicellular consortia where cellular strains communicate through chemical signals in liquid media or agar plates, resulting in stable configurations or patterns under controlled conditions. Programmable synthetic assemblies utilized differential adhesion models to drive self-organized cellular aggregates, achieving predictable spatial arrangements based on specific initial conditions. Synthetic morphology experiments, including the development of Xenobots and Anthrobots, emphasized the emergence of functionalities like memory and homeostasis without genetic editing, showcasing the balance between engineered predictability and emergent biological behaviors. The results highlight both the successes and inherent challenges in scaling synthetic multicellular systems, pointing to the necessity for novel methodologies to manage complexity and unpredictability in larger-scale constructs.
@@
Advancements in bioinformatics and network analysis have further elucidated the connections between bioelectric signaling and disease. Pio-Lopez and Levin [36] introduce MultiXVERSE, a universal multilayer network embedding method applied to investigate the causal link between GABA neurotransmitters and cancer. By constructing a multiplex-heterogeneous network that combines drug, disease, and gene interactions, MultiXVERSE employs Random Walks with Restart (RWR) for node similarity calculations and link prediction via a Random Forest classifier. The study involved creating a multilayer network incorporating protein-protein interactions, drug interactions from sources like DrugBank, and disease networks representing symptom similarity and comorbidity relationships. Experimental validation using Xenopus laevis embryos exposed to the GABA agonist muscimol demonstrated that fluoxetine delivery via a wearable bioelectronic bandage significantly enhanced wound healing, evidenced by a 39.9% increase in re-epithelialization and a 27.2% reduction in the M1/M2 macrophage ratio. These findings suggest that bioelectric states, modulated by neurotransmitters like GABA, play a significant role in tumorigenesis and cancer progression, offering potential avenues for therapeutic interventions through bioelectric modulation.
@@
Exploring the evolutionary aspects of multicellular intelligence, Hartl, Risi, and Levin [14] investigate how different levels of cellular competency within a multi-scale competency architecture (MCA) affect the evolutionary process. Their in-silico model simulates morphogenesis using neural cellular automata (NCAs), where cellular agents evolve to form target patterns. The methodology involves setting up NCAs on an 8x8 grid to form predefined patterns, such as a Czech flag, through coordinated cell state updates. Each cell contains an artificial neural network (ANN) that processes information from neighboring cells and proposes state changes. Key competency parameters, including decision-making probability (PD) and redundancy in decision pathways (R), are adjusted to simulate various levels of cellular reliability and computational capacity. The evolutionary fitness of each NCA is measured based on the accuracy and stability of forming the target pattern, with cells experiencing noise to simulate environmental challenges. Results indicate that higher competency levels accelerate evolutionary success and enhance the robustness and adaptability of morphogenetic processes. Additionally, NCAs with multi-scale competency encoding demonstrated better generalization to new environmental conditions compared to direct encoding models, suggesting that cellular-level intelligence significantly contributes to the adaptability and resilience of developing structures.
@@
The concept of memory as a dynamic and adaptive process is further explored by Levin [19], who posits that biological memories are not static repositories but are continuously reinterpreted to maintain functional relevance during significant physiological changes. Through conceptual analysis and comparative studies, Levin illustrates how memory engrams adapt to new biological contexts, such as metamorphosis in caterpillars and planarian regeneration. For instance, during metamorphosis, memories associated with a caterpillar’s sensory-motor systems are reformatted to remain functionally relevant for the butterfly’s needs, despite drastic physical and behavioral transformations. Similarly, planaria can transfer memories through body regeneration, revealing that specific memory cues adapt across different biological states. This phenomenon, termed "mnemonic improvisation," is central to understanding intelligence, as memory supports flexible, real-time problem-solving and continuity of self. Levin’s perspective offers a framework for understanding intelligence as an emergent property of dynamic memory processes that evolve across different biological scales.
@@
Investigating the emergence of higher-level agency, Tissot et al. [45] examine how phase synchronization among individual decision cycles in a decentralized system can lead to the emergence of higher-level agency. Their computational model simulates collective behaviors of individual units that engage in synchronized decision-making, achieving forms of collective action that surpass the capabilities of individual units acting alone. Each unit operates within a modular system, interacting more frequently with intra-module units than with inter-module units. The decision cycles are governed by oscillatory functions where phase adjustments occur based on the alignment of decisions within each module. Results demonstrate that synchronization leads to spontaneous collective shifts from local to global optima, overcoming individual-level stress barriers. After synchronization, modules coordinate responses across the entire system, achieving states with minimized collective stress. These findings suggest that synchronization mechanisms are fundamental to the emergence of collective intelligence in decentralized biological systems, enabling higher-level problem-solving competencies.
@@
In cancer research, Kofman and Levin [16] perform a meta-analysis evaluating pharmacological agents targeting ion channels as potential cancer therapies. Their comprehensive literature search across databases like IUPHAR, DrugBank, and PubMed identified various ion channel-targeting drugs, categorizing them based on their cancer-promoting or inhibiting effects. The analysis revealed that numerous ion channel drugs exhibit significant cancer-modifying properties across multiple cancer types, including breast, liver, brain, and ovarian cancers. Voltage-gated sodium and potassium channels frequently appeared in both pro- and anti-cancer categories, suggesting that the bioelectric state of cells can significantly influence tumor progression and metastasis. Some drugs demonstrated dual effects, indicating that factors such as timing, dosage, or biological context might influence their impact on cancer phenotypes. These findings highlight the therapeutic potential of bioelectric modulation in cancer treatment strategies, offering new targets for intervention beyond traditional genetic approaches.
@@
Zhang, Goldstein, and Levin [49] explore morphogenetic processes through the application of classical sorting algorithms. By implementing sorting algorithms such as Bubble Sort, Insertion Sort, and Selection Sort in a decentralized, cell-view model, they simulate self-sorting cellular arrays capable of handling error-prone conditions through emergent behaviors like clustering and delayed gratification. The methodology involved treating each element in an array as an autonomous agent capable of viewing its neighboring elements and deciding whether to swap based on local conditions. In addition to normal, fully-functional cells, the study introduced "frozen" cells that could either initiate no swaps or be immovable by others, simulating unreliability as seen in biological media. Results demonstrated that cell-view algorithms maintained sorting efficiency and robustness even in the presence of unreliable cells, showcasing emergent problem-solving behaviors in minimalistic systems. These findings indicate that algorithmic models can effectively mimic biological self-organization and problem-solving capabilities, offering insights into the decentralized control mechanisms inherent in biological morphogenesis.
@@
Theoretical frameworks for self-organization during prenatal development are advanced by Ciaunica et al. [7], who apply the concept of Markov blankets to model the maternal-fetal relationship during human prenatal development. Their study emphasizes the interplay between maternal and fetal immune systems, conceptualizing pregnancy as a dynamic process where two immune systems operate as self-organizing entities with distinct yet interconnected Markov blankets. The placenta serves as a boundary regulating nutrient and signal exchange, maintaining separate yet coordinated environments for both the mother and fetus. The methodology involved a detailed analysis of immune interactions, including the migration of immune cells like macrophages and natural killer cells through the placenta, contributing to the unique immune environment of pregnancy. The results highlight a theoretical framework where the Markov blanket formalism provides insights into the maternal-fetal relationship as an instance of nested selves, facilitating co-homeostasis and coordinated developmental processes. This perspective offers a novel understanding of biological self-organization, emphasizing the role of shared Markov blankets in maintaining the balance and stability of interconnected yet distinct biological systems during prenatal development.
@@
Decentralized control mechanisms in microswimmers are investigated by Hartl, Levin, and Zöttl [13], who demonstrate that artificial neural networks (ANNs) and genetic algorithms can optimize collective locomotion strategies in N-bead swimmers modeled after the Najafi-Golestanian (NG) microswimmer. The methodology involved constructing a virtual N-bead swimmer model where each bead acts as an independent agent capable of perceiving the states of its neighboring beads and applying forces accordingly. Using reinforcement learning techniques, the ANN parameters were evolved to maximize the swimmer’s speed, with the study testing swimmer sizes ranging from three to 100 beads. Results showed that the decentralized approach enabled robust and efficient swimming gaits, with locomotion strategies scaling effectively as the number of beads increased. Type B swimmers, in particular, displayed collective contractions reminiscent of caterpillar-like movement. Additionally, the evolved locomotion strategies demonstrated resilience to morphological changes and adaptability to different swimmer configurations without retraining, highlighting the potential of decentralized systems in achieving coordinated movement in autonomous microswimmers for biomedical applications.
@@
Levin [22] theorizes that bioelectric networks in multicellular systems establish cognitive boundaries that enable cells to function as unified cognitive entities. By integrating developmental biology, bioelectric signaling, and cognitive science, Levin posits that bioelectric gradients serve as computational mediums guiding multicellular coordination and pattern formation. The methodology involves a theoretical framework combining bioelectric signaling mechanisms, such as ion channels and gap junctions, with cognitive boundary concepts to model how cells collectively process information and regulate behavior. Examples from studies on Xenopus embryos, planarians, and tadpole regeneration support the hypothesis that bioelectric states function as pre-patterns, directing cellular organization and overriding genetic instructions to achieve specific anatomical outcomes. The results indicate that bioelectric signaling underpins cognitive-like processes across biological scales, suggesting that the "self" emerges from bioelectric boundaries that define the spatio-temporal limits within which cells interact and maintain coherence as a unified system.
@@
Lyon et al. [25] expand the understanding of cognitive processes in biology by redefining basal cognition to include sensory and information-processing mechanisms observed across diverse life forms. Their study examines cognitive behaviors in single-celled organisms, prokaryotes, and plants, focusing on capacities such as sensing, perception, memory, decision-making, and communication. By investigating mechanisms like quorum sensing in bacteria, cellular adhesion in amoebas, and bioelectrical signaling in plants, the authors argue that foundational cognitive capacities are evolutionarily conserved and precede the development of nervous systems. Experiments involved monitoring cellular responses to nutrient gradients and light stimuli, eliciting directional movement and adaptive behaviors without the presence of a nervous system. The results reveal that even simple organisms exhibit basic cognitive functions, suggesting a continuity of cognitive processes across biological complexity. This foundational understanding informs the study of more advanced cognitive systems and has potential applications in regenerative medicine and artificial intelligence, where principles of basal cognition can inspire novel approaches to problem-solving and adaptive behaviors.
@@
In "Resting Potential, Oncogene-induced Tumorigenesis, and Metastasis: The Bioelectric Basis of Cancer in vivo," Lobikin et al. [24] utilize Xenopus laevis tadpoles to investigate how bioelectric signaling influences cancerous transformations. The methodology involved modulating transmembrane voltage potentials (Vmem) by depolarizing specific GlyCl-expressing instructor cells using pharmacological agents like ivermectin to open glycine-gated chloride channels. By manipulating extracellular chloride concentrations, controlled depolarization was induced, leading to melanocytes exhibiting metastatic-like transformations, including hyperproliferation, morphological changes, and invasive behavior. Additionally, exposure to the carcinogen 4-Nitroquinoline 1-oxide (4NQO) generated localized tumors, with bioelectric changes visualized using fluorescent voltage- and ion-reporter dyes. The study found that oncogene-induced tumors displayed elevated sodium levels, potentially useful for non-invasive diagnostics. Conversely, forced hyperpolarization of cells via expression of hyperpolarizing ion channels mitigated oncogene-induced tumorigenesis, demonstrating that bioelectric state modulation can suppress tumor formation. These findings highlight the active role of bioelectric signaling in cancer progression and metastasis, suggesting bioelectric modulation as a promising strategy for cancer therapeutics.